OpenStudy (anonymous):

Find the constant of variation k for the direction variation 2x + 6y = 0

hartnn (hartnn):

isolate y subtract 2x from both sides

OpenStudy (anonymous):

oh, alright. I wasn't sure if I should still isolate y since now there was two numbers/two letters..Okay, so that'd be 4y = 0? O.o

hartnn (hartnn):

not actually, you combine only like terms :) 2x and 6y are not like terms.

OpenStudy (anonymous):

but then it'd be 6y = -2..O.o

hartnn (hartnn):

let me give you a similar example : 10x+3y = 0 to isolate y, we observe that 10x is ADDED to the y term so we SUBTRACT 10x from both sides, 10x -10x +3y = 0 -10x which gives 3y = -10x ---------------- now we see that 3 is MULTIPLIED with y so to isolate y, we DIVIDE both sides by 3 3y/3 = -10x/3 y = -10x/3 and comparing with y= kx k = -10/3

hartnn (hartnn):

if you meant \(\large 6y =-2x\) then you're absolutely correct :)

OpenStudy (anonymous):

Oh yeah, I forget to add the letters all the time XD Yeah..So then I'd..Divide 6 by both sides?

hartnn (hartnn):

correct! good, divide both sides by 6

OpenStudy (anonymous):

well wouldn't that be -3? O.o

OpenStudy (anonymous):

Oh wait wait wait

OpenStudy (anonymous):

That would be 0.3 repeating, right?

OpenStudy (anonymous):

And then I'd have to put the 0.3 repeating as a fraction, right?

OpenStudy (anonymous):

So then the answer would be -1/3? O.o

hartnn (hartnn):

6y =-2x 6y/6 = -2x/6 y = -2x/6 try to reduce -2/6 answer in fractions is allowed, right ?

hartnn (hartnn):

k =-1/3 is correct :)

OpenStudy (anonymous):

Yay :D Tysm.

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

I also got -1/3x for my answer for this problem..Is that correct? I got it by basically doing the same thing as I did here.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

y =mx -2 = 6m m = -1/3 y =-1/3x yes, you are correct! :) good job!